# Residual Neural Terminal Constraint for MPC-based Collision Avoidance in Dynamic Environments

### v1

@@ Abstract
In this paper, we propose a hybrid MPC local planner that uses a learning-based approximation of a time-varying safe set, derived from local observations and applied as the MPC terminal constraint. This set can be represented as a zero-superlevel set of the value function computed via Hamilton-Jacobi (HJ) reachability analysis, which is infeasible in real-time. We exploit the property that the HJ value function can be expressed as a difference of the corresponding signed distance function (SDF) and a non-negative residual function. The residual component is modeled as a neural network with non-negative output and subtracted from the computed SDF, resulting in a real-time value function estimate that is at least as safe as the SDF by design. Additionally, we parametrize the neural residual by a hypernetwork to improve real-time performance and generalization properties. The proposed method is compared with three state-of-the-art methods in simulations and hardware experiments, achieving up to 30% higher success rates compared to the best baseline while requiring a similar computational effort and producing high-quality (low travel-time) solutions.
@@ 1 Introduction
One of the most prominent techniques for trajectory planning and control in mobile robotics is model predictive control (MPC), which consists of a few core components such as a robot model, cost function, set of constraints, terminal set, and optimization algorithm [3]. Even though a proper choice of any component is not trivial, designing a terminal set that is control invariant is particularly difficult, especially in dynamic environments. This set is important to achieve the recursive feasibility of the MPC, which is a desired property in any safety-critical application.
To determine the maximal control invariant set, which we call maximal safe set, Hamilton-Jacobi (HJ) reachability analysis can be used. This method relies on the dynamic programming principle to solve for the HJ value function and therefore cannot be executed online. To overcome this issue, we propose a learning-based approach to approximate the value function based on local observations. Moreover, we exploit the property that the HJ value function is always less than or equal to the corresponding signed distance function (SDF). This allows us to estimate only the residual function using a neural model with non-negative output and subtract it from the numerically computed SDF. In this way, the learned HJ value function approximation is guaranteed, by design, to be at least as safe as the corresponding SDF. The main contributions of this paper are:
  • Learning-based estimation of maximal safe sets for dynamic environments in real-time, with a model architecture that guarantees, by design, the same or higher level of safety compared to an SDF-based approach.
  • Extensive empirical comparison of the proposed method with three state-of-the-art methods in simulations and on a physical robot.
@@ 2 Related Work
Robot navigation often requires some assumptions about the environment in which the robot is deployed. If the environment is shared with other agents, a standard differentiation is whether the agents are interactive or not. In the first case, the robot and other agents collaboratively avoid collisions, while in the second case, the robot treats other agents as non-interactive moving obstacles. Some classical algorithms for interactive planning, such as ORCA [4], are based on the velocity obstacle (VO) paradigm [5]. The ORCA method is also combined with an MPC planner to formulate velocity constraints [6], predict human motion [7], or refine diffusion-based motion prediction [8]. Another popular approach is to learn interaction-aware motion prediction and then integrate it within the MPC planner as an environment model [9].
As indicated in Section 1, our work is focused on non-interactive scenarios where the assumption about reciprocal collision avoidance might lead to an unsafe situation. Similar to the interactive case, methods used in such environments are usually based on distance functions [10] or define constraints in the velocity space [5]. Focusing on MPC-based planners, one can use SDF [11, 12] or VO [13] to define collision avoidance constraints. Another approach is introduced in [14], where the authors combine MPC with a discrete-time control barrier function (CBF). However, none of these methods explicitly consider terminal set design and rely on a sufficiently long prediction horizon for improved recursive feasibility. On the other hand, our aim is to design a control invariant set in real-time and employ it as the terminal set for the MPC planner.
The maximal control invariant set can be determined using HJ reachability analysis, which is a framework for model-based computation of reachable sets [15]. This framework suffers from the curse of dimensionality and is impractical for time-critical systems, which motivated methods such as efficient initialization, system decomposition, and neural approximation of the HJ value function [16, 17, 18]. However, those approaches assume complete a priori knowledge about the environment, which is often not satisfied, specifically for dynamic environments. The work in [19] introduced parameter-conditioning by augmenting the system dynamics with a set of “virtual” states, and the idea was applied in [20] and [21] to motion planning tasks. The main limitation of this method is that the dimensionality of the problem increases with the number of parameters. To overcome this, authors in [22] propose a learning-based HJ value function estimation based on high-dimensional environment observations. Yet, this method was applied only to static environments without any formal guarantees on the approximation error. In our work, we extend this idea to dynamic environments and propose a new model architecture that guarantees that the value function estimate is at least as safe as SDF.
@@ 3.1 Problem Formulation
In the formulation above, $x_{k+i}$ and $u_{k+i}$ are predicted state and control vectors at future time $k+i$ made at the current time step $k$ over the prediction horizon of length $N$. Constraint Eq. 3b enforces the system’s dynamics, Eq. 3c and Eq. 3d ensure that the predicted state and control vectors have admissible values, while Eq. 3e and Eq. 3f represent stage and terminal constraints, respectively. For the stage cost $\ell(x,u)$ and the terminal cost $\ell_{N}(x)$ we use the standard quadratic form:
@@ 3.2 Hamilton-Jacobi Reachability Analysis in Time-Varying Environments
The set $\mathcal{F}(t)$ is defined as the zero-sublevel set of a failure function $F(x,t)$ (SDF in our context), i.e. $\mathcal{F}(t)=\{x:F(x,t)\leq 0\}$}, and to compute BRT, we define cost function
$J(x,t,u(\cdot))=\min_{\tau\in[t,T]}F(\xi^{u}_{x,t}(\tau),\tau).$ (7)
Since $\mathcal{F}(t)$ represents occupied space at time $t$, the goal is to find the optimal control that will maximize cost $J$. Therefore, we introduce the value function
$V(x,t)=\sup_{u(\cdot)}\{J(x,t,u(\cdot))\}.$ (8)
The value function is computed over the system state space using the dynamic programming principle, resulting in the following Hamilton-Jacobi-Bellman Variational Inequality (HJB-VI) [24, 23]:
@@ 4 Methodology
In this section, we present our framework for online estimation of the maximal safe set in unknown dynamic environments. The goal is to estimate the maximal safe set in real-time and integrate it within the MPC local planner as the terminal constraint. As explained in Section 3.2, the maximal safe set at some time $t^{\prime}$ is the zero-superlevel set of the HJ value function $V(x,t^{\prime})$ and it depends on the evolution of the failure set $\mathcal{F}(t)$ for $t>t^{\prime}$. If we consider the MPC problem formulation defined for time $k\cdot\delta t$, we need to estimate the maximal safe set at time $(k+N)\cdot\delta t$}, i.e. the value function $V(x,(k+N)\cdot\delta t)$. To do so, it is necessary to reason about the evolution of the environment and the motion of the other agents even after the prediction horizon. This setting is very different from static environments considered in [22], where the failure set does not change over time. In this case, we need to introduce a certain assumption about the motion of the agents even after the final time step. In this paper, it is assumed that the moving obstacles will continue to move with constant velocity, but other assumptions can be adopted as well (see Section 7). This allows us to condition the safe set estimation at time $(k+N)\cdot\delta t$} based on predictions over the time horizon $k\cdot\delta t,(k+N)\cdot\delta t]$}.
To achieve real-time performance, we rely on the idea introduced in [22] where the model is separated into two neural networks — a hypernetwork and a main network. The hypernetwork is an expressive model with trainable parameters that parametrizes the main network based on the conditioning input. Complementary, the main network is a lightweight model that is parametrized by the hypernetwork and integrated within numerical optimization. In our approach, we use the tail of the SDF sequence corresponding to the predicted motion of the surrounding agents as the input to the hypernetwork as follows. At time step $k$, we obtain predicted positions $\{p_{k}^{j},p_{k+1}^{j},\ldots,p_{k+N}^{j}\}_{j=1}^{M}$} for $M$ agents over the prediction horizon and compute the corresponding SDF sequence $\{F_{k}(x),F_{k+1}(x),\ldots,F_{k+N}(x)\}$}. The first $N$ SDFs are used as the stage constraints, i.e. $h_{k+i}(x)=F_{k+i}(x)$ for $i=0,1,\ldots,N-1$, and the last $K$ SDFs, from NK+1N-K+1 to $N$, are used as the input to the hypernetwork to generate main network parameters $\Theta$.
@@
Figure 2: Training diagram of the proposed method. The training data is split into future and past parts of the observed trajectories. The past part is used to compute the SDF sequence, which is fed into the hypernetwork to parametrize the main network. The future part is used to compute the true HJ value function. The predicted value function is obtained as the difference between the numerically computed SDF at the current time and the estimated residual.
@@ 5 Experimental Results
Model Training To train the model, we first create a dataset of obstacle trajectories in a simulation environment using a constant velocity model [^1] The dataset of trajectories can be generated using any other motion model or by collecting real-world data.. We assume that the perception field of the robot is limited to a square area of 8x8 m centered at the robot’s position. Positions and velocity vectors of the local obstacles are randomly selected, and the trajectories are simulated forward in time. Next, as explained in Section 4, we split every trajectory into past and future segments and compute SDF sequences and the true value functions. As a hypernetwork, we use a deep CNN model that takes an SDF sequence as input and outputs the main network parameters. The main network is a simple MLP model with three inputs and a single output with $ELU+1$ activation function, which estimates residual over the state space grid. More details can be found in Appendix C.
@@
During the experiments, we consider standard metrics for mobile robot navigation. The first metric is success rate, i.e., the percentage of experiments completed without collision, which is presented in Fig. 3 (left). The results show a higher success rate of the proposed RNTC-MPC method compared to different baselines, regardless of the MPC horizon length. Fig. 3 (right) shows the mean optimization time for the evaluated MPC methods, illustrating that the proposed method achieves comparable real-time performance.
@@
Figure 4: (Left) Travel times during navigation between initial and goal positions. (Right) Trade-off between success rate and mean travel time. Our method achieves the best compromise between the two objectives.
@@
Hardware Experiments In addition to simulations, we assess the proposed method on a real robot and compare it to the main baselines: SDF-MPC, DCBF-MPC, and VO-MPC. We use a mobile robot equipped with 3D LiDAR and stereo cameras for our experiments. All computation and processing tasks are performed onboard. Similar to the simulation experiments, the task is to move from the initial position to the goal position while avoiding moving obstacles. We perform 10 trials for each method in scenarios where pedestrians intentionally intersect robot motion. The results are presented in Table 1, illustrating 30% higher success rate compared to the baselines, while achieving the lowest travel time. Additional details and qualitative results for hardware experiments are presented in Appendix E and the supplemental video. The discrepancy in results between simulations and hardware experiments can be explained by the lower number of trials in the hardware experiments.
@@ Appendix A Ablation Study on Model Architecture
Figure 7: Two samples of HJ value function slices for $\theta=0$ rad (right - full slice; left - zoomed part of the slice). The results illustrate that the value function approximation in the RNTC method classifies states as unsafe if they are unsafe according to the SDF, which is not true for the NTC method.
@@ Appendix B CME Loss Function Analysis
The main motivation for the proposed hybrid CME loss function is to achieve faster convergence during the training process and better final performance. We compare the proposed loss defined in Eq. 15 with the standard MSE loss function used in regression tasks. We train the same model under the same settings and on the same dataset. Fig. 8 (left) shows IoU metric on the validation dataset during training. The proposed CME loss converges significantly faster, and the final IoU value at the end of training is higher.
@@ Appendix C Model Training
To train the model, we generate a dataset of obstacle trajectories using the constant velocity model for obstacle motion. It is important to mention that the dataset of trajectories can be generated using any other motion model or by collecting real-world data. The trajectories are split at some point in time into past and future segments, which are then used to calculate corresponding SDF sequences and true HJ value functions. In this paper, we assume that the robot can sense moving obstacles within a square region of size 8x8 meters centered at the robot’s location. The positions and velocity vectors of the local obstacles are randomly sampled within this region, and the trajectories are then simulated forward in time. We assume that the maximal number of obstacles is 4 and that they can move in any direction with a maximal speed of up to 1 m/s. In total, we generate 40.000 training pairs, i.e., pairs of SDF sequences and the corresponding HJ value function.
For HJ value function computation, we use hj_reachability [^2] https://github.com/StanfordASL/hj_reachability Python library with modifications to handle time-varying failure sets. The value function is computed over a discrete state space grid of size 100x100x40. The computation is performed on a workstation with an NVIDIA RTX3090 GPU and the process took approximately 20h to complete.
The hypernetwork, the main network, and the training process are implemented using the PyTorch framework. The hypernetwork is implemented as a standard CNN model that takes as input an SDF sequence and outputs the main network parameters $\Theta$. In our concrete case, the SDF sequence contains only 2 SDFs (at the current time and 0.4 s into the past) because we assumed the constant velocity model, and two SDFs contain enough information to forecast the future motion. The complete hypernetwork architecture is provided in LABEL:lst:hypernet_summary, resulting in $\sim 9.37\cdot 10^{6}$ trainable parameters. The main network is implemented as an MLP model with three input neurons (size of the state vector) and a single output with $ELU+1$ activation function to guarantee a non-negative estimated residual. The details are presented in LABEL:lst:main_net_summary, and the total number of parameters is 4519. The residual is subtracted for the SDF at the current time to estimate the HJ value function that corresponds to the input SDF sequence. Then the predicted value function is compared to the true value function via the loss function Eq. 15 (for $\gamma=0.2$, selected via grid search) and the weights of the hypernetwork are updated using mini-batch gradient descent. We train the model for 100 epochs using Adam optimizer with an initial learning rate of 0.0001 and a batch size of 40. The learning rate is reduced by a factor of 10 at epochs 85 and 95. The training process takes approximately 15h on an NVIDIA RTX3090 GPU.

### v2

@@ Abstract
In this paper, we propose a hybrid MPC local planner with a learning-based approximation of a time-varying safe set, which is derived from local observations and applied as the terminal constraint. This set can be represented as the zero-superlevel set of the value function computed via Hamilton-Jacobi (HJ) reachability analysis, which is often infeasible in real-time. We exploit the property that the HJ value function can be expressed as a difference of the corresponding signed distance function (SDF) and a non-negative residual function. The residual component is modeled as a neural network with non-negative output and subtracted from the computed SDF, resulting in a real-time value function estimate that is at least as safe as the SDF by design. Additionally, we parametrize the neural residual by a hypernetwork to improve real-time performance and generalization properties. The proposed method is compared with three state-of-the-art methods in simulations and hardware experiments, achieving up to 30% higher success rates compared to the best baseline while requiring a similar computational effort and producing high-quality (low travel-time) solutions.
@@ 1 Introduction
One of the most prominent techniques for motion planning and control in mobile robotics is model predictive control (MPC), which consists of a few core components such as a robot model, cost function, constraints, terminal set, and optimization algorithm [3]. Even though a proper choice of any component is not trivial, designing a terminal set that is control invariant is particularly difficult, especially in dynamic environments. This set is necessary to achieve the recursive feasibility of the MPC, which is a crucial property in safety-critical applications.
To determine the maximal control invariant set, which we call maximal safe set, Hamilton-Jacobi (HJ) reachability analysis can be used. This method relies on the dynamic programming principle to solve for the HJ value function and often cannot be executed online. To overcome this issue, we propose a learning-based approach to approximate the value function based on local observations. Moreover, we exploit the property that the HJ value function is always less than or equal to the corresponding signed distance function (SDF). This allows us to estimate only the residual function using a neural model with non-negative output and subtract it from the numerically computed SDF. In this way, the learned HJ value function approximation is guaranteed, by design, to be at least as safe as the corresponding SDF. The main contributions of this paper are:
  • A learning-based method for real-time estimation of the maximal safe set in dynamic environments with a model architecture that guarantees, by design, the same or higher level of safety compared to an SDF-based approach.
  • Extensive empirical comparison of the proposed method with state-of-the-art methods in simulations and on a physical robot.
@@ 2 Related Work
Robot navigation often requires some assumptions about the environment in which the robot is deployed. If the environment is shared with other agents, a standard distinction is whether the agents are interactive or not. In the first case, the robot and other agents collaboratively avoid collisions, while in the second case, the robot treats other agents as non-interactive moving obstacles. Some classical algorithms for interactive planning, such as ORCA [4], are based on the velocity obstacle (VO) paradigm [5]. The ORCA method is also combined with an MPC planner to formulate velocity constraints [6], predict human motion [7], or refine diffusion-based motion prediction [8]. Another popular approach is to learn interaction-aware motion prediction and then integrate it within the MPC planner as an environment model [9].
As indicated in Section 1, our work is focused on non-interactive scenarios where the assumption about reciprocal collision avoidance might lead to an unsafe situation. Similar to the interactive case, methods used in such environments are usually based on distance functions [10] or velocity obstacles [5]. Focusing on MPC-based planners, one can use SDF [11] or VO [12] to define collision avoidance constraints. Another approach is introduced in [13], where the authors combine MPC with a discrete-time control barrier function (CBF). However, none of these methods explicitly consider terminal set design and rely on a sufficiently long prediction horizon for improved recursive feasibility. On the other hand, our aim is to design a control invariant set in real-time and employ it as the terminal set for the MPC planner.
The maximal control invariant set can be determined using HJ reachability analysis, which is a framework for model-based computation of reachable sets [14]. This framework suffers from the curse of dimensionality and is impractical for time-critical systems, which motivated methods such as efficient initialization, system decomposition, and neural approximation of the HJ value function [15, 16, 17]. However, those approaches assume complete a priori knowledge about the environment, which is often not satisfied, especially for dynamic environments. The work in [18] introduced parameter-conditioning by augmenting the system dynamics with a set of “virtual” states, and the idea was applied in [19] and [20] to motion planning tasks. The main limitation of this method is that the dimensionality of the problem increases with the number of parameters. To overcome this, authors in [21] propose a learning-based HJ value function estimation based on high-dimensional environment observations. Yet, this method was applied only to static environments without any formal guarantees on the approximation error. In our work, we extend this idea to dynamic environments and propose a new model architecture that guarantees that the value function estimate is at least as safe as the SDF.
@@ 3.1 Problem Formulation
In the formulation above, $x_{k+i}$ and $u_{k+i}$ are predicted state and control vectors at future time $k+i$ made at the current time step $k$ over the prediction horizon of length $N$. Constraint Eq. 3b enforces the system’s dynamics, Eq. 3c and Eq. 3d ensure that the predicted state and control vectors have admissible values, while Eq. 3e and Eq. 3f represent stage and terminal constraints, respectively. For the stage cost $\ell(x,u)$ and terminal cost $\ell_{N}(x)$ we use the standard quadratic form:
@@ 3.2 Hamilton-Jacobi Reachability Analysis in Time-Varying Environments
The set $\mathcal{F}(t)$ is defined as the zero-sublevel set of a failure function $F(x,t)$ (SDF in our context), i.e. $\mathcal{F}(t)=\{x:F(x,t)\leq 0\}$}, and to compute the BRT, we define cost function
$J(x,u(\cdot),t)=\min_{\tau\in[t,T]}F(\xi^{u}_{x,t}(\tau),\tau).$ (7)
Since $\mathcal{F}(t)$ represents occupied space at time $t$, the goal is to find the optimal control that will maximize the cost $J$. Therefore, we introduce the value function
$V(x,t)=\sup_{u(\cdot)}\{J(x,u(\cdot),t)\}.$ (8)
The value function is computed over the system state space by solving the following Hamilton-Jacobi-Bellman Variational Inequality (HJB-VI) using the dynamic programming principle [23, 22]:
@@ 4 Methodology
In this section, we present our framework for online estimation of the maximal safe set in unknown dynamic environments. The goal is to estimate this set in real-time and integrate it within the MPC local planner as the terminal constraint. As explained in Section 3.2, the maximal safe set at some time $t^{\prime}$ is the zero-superlevel set of the HJ value function $V(x,t)$ and it depends on the evolution of the failure set $\mathcal{F}(\tau)$ for $\tau>t$. If we consider the MPC problem formulation defined for time $k\cdot\delta t$, we need to estimate the maximal safe set at time $(k+N)\cdot\delta t$}, i.e. the value function $V(x,(k+N)\cdot\delta t)$. To do so, it is necessary to reason about the evolution of the environment and the motion of the other agents even after the prediction horizon. This setting is very different from static environments considered in [21], where the failure set does not change over time. In this case, we need to introduce a certain assumption about the motion of the agents even after the final time step. In this paper, it is assumed that the moving obstacles will continue to move with constant velocity, but other assumptions can be adopted as well (see Section 7). This allows us to condition the safe set estimation at time $(k+N)\cdot\delta t$} based on predictions over the time horizon $k\cdot\delta t,(k+N)\cdot\delta t]$}.
To achieve real-time performance, we rely on the idea introduced in [21] where the model is separated into two neural networks — a hypernetwork and a main network. The hypernetwork is an expressive model with trainable parameters that parametrizes the main network based on the conditioning input. Complementary, the main network is a lightweight model that is parametrized by the hypernetwork and integrated within numerical optimization as a constraint function. In our approach, we use the tail of the SDF sequence, corresponding to the predicted motion of the surrounding agents, as the input to the hypernetwork. At time step $k$, we obtain predicted positions $\{p_{k}^{j},p_{k+1}^{j},\ldots,p_{k+N}^{j}\}_{j=1}^{M}$} for $M$ agents over the prediction horizon and compute the corresponding SDF sequence $\{F_{k}(x),F_{k+1}(x),\ldots,F_{k+N}(x)\}$}. The first $N$ SDFs are used as the stage constraints, i.e. $h_{k+i}(x)=F_{k+i}(x)$ for $i=0,1,\ldots,N-1$, and the last $K$ SDFs, from NK+1N-K+1 to $N$, are used as the input to the hypernetwork to generate main network parameters $\Theta$.
@@
Figure 2: Training diagram of the proposed method. The training data is split into future and past parts of the observed trajectories. The past part is used to compute the SDF sequence, which is fed into the hypernetwork to parametrize the main network. The future part is used to compute the true HJ value function. The predicted value function is obtained as the difference between the numerically computed SDF at the current time and the neural residual estimate.
@@ 5 Experimental Results
Model Training To train the model, we first create a dataset of obstacle trajectories in a simulation environment using a constant velocity model. We assume that the perception field of the robot is limited to a square area of $\times8$ m centered at the robot’s position. Position and velocity vectors of the local obstacles are randomly selected, and the trajectories are simulated forward in time. Next, as explained in Section 4, we split every trajectory into past and future segments and compute SDF sequences and the true value functions. As a hypernetwork, we use a deep CNN model that takes an SDF sequence as input and outputs the main network parameters. The main network is a simple MLP model with three inputs and a single output with $ELU+1$ activation function, which estimates residual over the state space grid. More details can be found in Appendix C.
@@
During the experiments, we consider standard metrics for mobile robot navigation. The first metric is success rate, i.e., the percentage of experiments completed without collisions, which is presented in Fig. 3 (left). The results show a higher success rate of the proposed RNTC-MPC method compared to different baselines, especially for short MPC horizon lengths. Fig. 3 (right) shows the mean optimization time for the MPC methods, illustrating that the proposed method achieves comparable real-time performance.
@@
Figure 4: (Left) Mean travel times during navigation between initial and goal positions. (Right) Trade-off between success rate and mean travel time. Our method achieves the best compromise between the two objectives.
@@
Hardware Experiments In addition to simulations, we assess the proposed method on a real robot and compare it to the main baselines: SDF-MPC, DCBF-MPC, and VO-MPC. We use a mobile robot equipped with 3D LiDAR and stereo cameras for our experiments. All computation and data processing is performed onboard. Similar to the simulation experiments, the task is to move from the initial position to the goal position while avoiding moving obstacles. We perform 10 trials for each method in scenarios where pedestrians intentionally intersect robot motion. The results are presented in Table 1, illustrating 30% higher success rate compared to the baselines, while achieving the lowest travel time. Additional details and qualitative results for hardware experiments are presented in Appendix E and the supplemental video [^1] https://www.youtube.com/watch?v=eyaAYIzgBpI. The discrepancy in results between simulations and hardware experiments can be explained by the lower number of trials in the hardware experiments.
@@ Appendix A Ablation Study on Model Architecture
Figure 7: Two samples of HJ value function slices for $\theta=0$ rad (right - full slice; left - zoomed slice). The results illustrate that the value function approximation in the RNTC method classifies states as unsafe if they are unsafe according to the SDF, which is not true for the NTC method.
@@ Appendix B CME Loss Function Analysis
The main motivation for the proposed hybrid CME loss function is to achieve faster convergence during the training process and better final performance. We compare the proposed loss defined in Eq. 15 with the standard MSE loss function used in regression tasks. We train the same model under the same settings and on the same dataset. Fig. 8 (left) shows IoU metric on the validation dataset during training. The proposed CME loss converges significantly faster, and the final IoU value at the end of training is higher [^2] To prevent numerical issues in the first epoch caused by the exponential term in the CME loss, one can pretrain the model using the MSE loss for a single epoch.
@@ Appendix C Model Training
To train the model, we generate a dataset of obstacle trajectories using the constant velocity model for obstacle motion. It is important to mention that the dataset of trajectories can be generated using any other motion model or by collecting real-world data. The trajectories are split at some point in time into past and future segments, which are then used to calculate corresponding SDF sequences and true HJ value functions. In this paper, we assume that the robot can sense moving obstacles within a square region of size $\times8$ meters centered at the robot’s location. The positions and velocity vectors of the local obstacles are randomly sampled within this region, and the trajectories are then simulated forward in time. We assume that the maximal number of obstacles is 4 and that they can move in any direction with a maximal speed of up to 1 m/s. In total, we generate 40.000 training pairs, i.e., pairs of SDF sequences and the corresponding HJ value function.
For HJ value function computation, we use the hj_reachability [^3] https://github.com/StanfordASL/hj_reachability Python library with modifications to handle time-varying failure sets. The value function is computed over a discrete state space grid of size $\times100$$\times$30. The computation is performed on a workstation with two NVIDIA RTX3090 GPUs and the process took approximately 20h to complete.
The hypernetwork, the main network, and the training process are implemented using the PyTorch framework. The hypernetwork is implemented as a standard CNN model that takes as input an SDF sequence and outputs the main network parameters $\Theta$. In our concrete case, the SDF sequence contains only 2 SDFs (at the current time and 0.4 s into the past) because we assumed the constant velocity model, and two SDFs contain enough information to forecast the future motion. The complete hypernetwork architecture is provided in LABEL:lst:hypernet_summary, resulting in $\sim 9.37\cdot 10^{6}$ trainable parameters. The main network is implemented as an MLP model with three input neurons (size of the state vector) and a single output with $ELU+1$ activation function to guarantee a non-negative estimated residual. The details are presented in LABEL:lst:main_net_summary, and the total number of parameters is 4519. The residual is subtracted for the SDF at the current time to estimate the HJ value function that corresponds to the input SDF sequence. Then, the predicted value function is compared to the true value function via the loss function Eq. 15 (for $\gamma=0.1$, selected via grid search) and the weights of the hypernetwork are updated using mini-batch gradient descent. We train the model for 100 epochs using Adam optimizer with an initial learning rate of 0.0001 and a batch size of 40. The learning rate is reduced by a factor of 10 at epochs 85 and 95. The training process takes approximately 15h on an NVIDIA RTX3090 GPU.